OpenStudy (anonymous):

Rewrite the expression using radical exponents. (simplify completly) Sqrt(x^4+y^4)

OpenStudy (anonymous):

nothing you can do with \[\sqrt{x^4+y^4}\] unless you want to write it as \[\left(x^4+y^4\right)^{\frac{1}{2}}\]

OpenStudy (anonymous):

Thanks